# Fractional Brownian motion with mean-density interaction: a myopic self-avoiding fractional stochastic process

### v2

@@ II.2 Mean-density interaction
(where $\delta_{ij}$is the Kronecker delta) and the force term
@@
$=$ $-A\left.\frac{\partial}{\partial x}P_{\mathrm{int}}(x,t_{n})\right|_{x=x_{n}^{(j)}}$ (9)
[…]
@@ II.3 Simulation details
To implement the mean-density interaction, we consider ensembles of up to $N=128$ particles. The mean integrated density $P_{\mathrm{int}}(x,t_{n})$ is collected as a histogram with a narrow bin width $\Delta x=0.1$. To achieve a smooth mean integrated density for our moderately large ensemble sizes, we replace the $\delta$ functions in the definition (5) of $P_{\mathrm{int}}(x,t_{n})$ by Gaussians of variance 0.25. We have confirmed that small changes of the histogram bin width and the variance of the smoothing Gaussian only lead to minuscule changes of the results  [^2] If the $\delta$ functions in the definition (5) of $P_{\mathrm{int}}(x,t_{n})$ are not smoothed, however, the histogram of the mean integrated density $P_{\mathrm{int}}(x,t_{n})$ develops an unphysical oscillatory instability.. The gradient in the definition of the force (9) is computed via a simple two-point formula directly from the histogram. To ensure the robustness of the results, we have also performed test calculations using a higher-order (4-point) gradient formula as well as calculations that do not use a histogram at all but analytically compute the derivative of the sum of the accumulated Gaussians. The latter method is restricted to times $t\lessapprox 10^{4}$ because of the high numerical effort of keeping track of all the Gaussians. All gradient algorithms give the same results (within our error bars) for the available simulation times.
@@ III Results: Mean-squared displacement
Figure 2: Mean-squared displacement $\langle x^{2}\rangle$ of FBM with mean-density interaction vs. time $t$ for $\alpha=0.7$ and several values of the interaction strength $A$. The data are averages over 60 ensembles of 128 random walkers each. The resulting statistical errors are much smaller than the symbol size. The dashed line represents a fit of the long-time behavior for $A=2/5$ with $\langle x^{2}\rangle=c\,t^{4/3}$ whereas the dash-dotted line shows the FBM relation $\langle x^{2}\rangle=\sigma^{2}t^{0.7}$. Inset: crossover time $\tau_{x}$ from FBM to interaction-dominated behavior vs. interaction strength $A$. The solid line is a power law fit, yielding an exponent of 1.05-1.05- 1.05.
@@
The trajectories of non-interacting FBM with $\alpha=0.7$ shown in panel (b) are highly jittery (antipersistent) because of the negative correlations of subdiffusive ($\alpha<1$) fractional Gaussian noise. This jittery motion is still visible in the presence of the mean-density interaction shown in panel (a). However, it is superposed onto a more regular motion away from the starting point (clearly visible for the trajectories furthest away from the origin (in the tails of the probability density). The comparison with panel (c) demonstrates that the trajectories of FBM with mean-density interaction do not resemble those of non-interacting FBM for $\alpha=4/3$ even though the mean-squared displacements of both processes evolve as $\langle x^{2}\rangle\sim t^{4/3}$. The $\alpha=4/3$ FBM trajectories are less jittery, even on short time scales because the fractional Gaussian noise for $\alpha>1$ is positively correlated.
@@ IV Scaling theory
If we assume, on the other hand, that the length scale $b(t)$ increases more slowly than $t^{2/3}$, the hypothetical contribution of the forces to the displacement would increase faster than $t^{2/3}$, leading to a contradiction. The scaling theory therefore predicts that one-dimensional FBM with mean-density interaction in free space is dominated by the fractional Gaussian noise (and behaves like regular FBM) for $\alpha>4/3$ (i.e., $\gamma<2/3)$, whereas it is interaction-dominated for $\alpha<4/3$ (i.e., $\gamma>2/3$). This yields the following mean-squared displacement behaviors,
$\langle x^{2}\rangle\sim\left\{\begin{array}[]{ll}t^{4/3}&\qquad\textrm{for }\alpha<4/3\\ t^{\alpha}=t^{2-\gamma}&\qquad\textrm{for }\alpha>4/3\end{array}\right.~.$ (19)
@@
$\tau_{x}\sim\left(\frac{A}{\sigma}\right)^{2/(3\alpha-4)}$ (20)
This relation holds in the interaction-dominated regime, $\alpha<4/3$. For the case of $\alpha=0.7$, the exponent in the power law (20) evaluates to 1.053-1.053- 1.053, in excellent agreement with the data in the inset of Fig. 2.
@@ V Results: Probability densities
where $\Gamma$ is the upper incomplete Gamma function (for details, see Appendix B). The function (24) has a maximum (with a cusp) at $x=0$ and a Gaussian tail (up to a power-law prefactor) for large $x$.
@@
$P(x,t)=\frac{1}{b(t)}Z\left[\frac{x}{b(t)}\right]$ (25)
@@
[…]
@@
[…]
@@
Figure 8: Mean-squared displacement $\langle x^{2}\rangle$ of FBM with mean-density interaction vs. time $t$ for $\alpha=0.7$, interaction strength $A=1/40$ and different particle numbers $N$. The data are averages over 2048 ensembles ($N=1$), 1024 ensembles ($N=2$), 128 ensembles ($N=16$), and 16 ensembles ($N=128$). The dashed lines are fits of the long-time behavior for $N=1$ and $N=128$ with $\langle x^{2}\rangle=c\,t^{4/3}$.
@@
Figure 9: Integrated density $P_{\mathrm{int}}(x,t)$ at time $t=2^{20}$ for $\alpha=0.7$, interaction strength $A=1/40$ and different particle numbers $N$. The data are averages over 4096 ensembles ($N=1$), 2048 ensembles ($N=2$), 256 ensembles ($N=16$), and 32 ensembles ($N=128$).
@@ VII Generalizations
Repeating the arguments at the end of Sec.  IV, we conclude that the motion of FBM with mean-density interaction will be interaction dominated if the FBM anomalous diffusion exponent $\alpha$ is smaller than $2\delta$. In this case the mean-squared displacement is expected to behave as $\langle\mathbf{x}^{2}\rangle\sim t^{2\delta}$. If on the other hand, $\alpha>2\delta$, the motion will be dominated by the fractional Gaussian noise leading to $\langle\mathbf{x}^{2}\rangle\sim t^{\alpha}$. Equation (28) shows that $\delta$ decreases with increasing space dimensionality $d$. This implies that the marginal value of $\alpha$, below which the interactions dominate over the noise, decreases with increasing $d$. The result that the interaction effects are strongest in one dimension and decrease with increasing $d$ is perhaps not unexpected as crowding is more easily achieved in lower dimensions.
Let us specifically consider the cases $d=2$ and $d=3$ for linear gradient forces, $\lambda=1$. In two dimensions, eq. (28) reduces to $\delta=1/2$. This implies that $\alpha=2\delta=1$ is the marginal value of $\alpha$ that separates the noise-dominated and interaction-dominated regimes in two dimensions. We note that this agrees with the finding that $d=2$ is the critical dimension of the so-called “true” or myopic self-avoiding random walk [38] discussed in Sec.  VIII. Analogously, for $d=3$, we find $\delta=2/5$ from equation (28), implying a critical $\alpha=0.8$.
@@ Appendix A Interaction with finite decay time
$P_{\mathrm{tot}}(x,t_{n})$ $=$ $\sum_{m=1}^{n}e^{-(t_{n}-t_{m})/\tau_{d}}\,\sum_{j=1}^{N}\delta[x-x^{(j)}_{m}]~,\quad$ (29)
@@
To ensure that the three time regimes predicted by the scaling theory (see Table 1) are well separated, we have chosen a comparatively strong interaction, $A=1/5$, for these simulations. This reduces the crossover time $\tau_{x}$ from the initial FBM regime to the interaction-dominated regime. The data for $\alpha=0.4$ (shown in the upper panel of the figure) clearly feature the three predicted regimes. Initially, for time $t<\tau_{x}\approx 10$, the mean-squared displacement follows the FBM relation $\langle x^{2}\rangle\sim t^{0.4}$. At $\tau_{x}$, it crosses over to the interaction-dominated behavior $\langle x^{2}\rangle\sim t^{4/3}$ established in the main part of the paper. After the second crossover at $t=\tau_{d}$, the long-time behavior reaches the second interaction-dominated regime with $\langle x^{2}\rangle\sim t^{2/3}$. This can be seen clearly in the figure where the solid line for $\tau_{d}=10^{3}$ represents a power-law fit with the exponent 2/32/32 / 3.
@@ Appendix B Integrated density of FBM
[…]
@@
[…]
The integral over $z$ yields the upper incomplete Gamma function, which concludes the derivation of eq. (24). It is worth emphasizing that expression (24) fulfills the scaling form (11) with $b(t)=x_{\mathrm{rms}}(t)=\sigma t^{\alpha/2}$. This can be seen explicitly by rewriting (24) as
$P_{\mathrm{int}}(x,t)$ $=$ $\frac{t}{\sigma t^{\alpha/2}}\,Y\left(\frac{x}{\sigma t^{\alpha/2}}\right)$ (34)

### v3

@@ II.2 Mean-density interaction
(where $\delta_{ij}$ is the Kronecker $\delta$) and the force term
@@
$=$ $-A\left.\frac{\partial}{\partial x}P_{\mathrm{int}}(x,t_{n})\right|_{x=x_{n}^{(j)}}~.$ (9)
Here, $P_{\mathrm{int}}=P_{\mathrm{tot}}/N$ is the mean integrated density of the ensemble. The factor $1/N$ in the relation between the total integrated density $P_{\mathrm{tot}}$ and the force is introduced in the spirit of mean-field theory to permit a well-defined thermodynamic limit $N\to\infty$. The parameter $A$ controls the character and strength of the interaction. For positive $A$, the particles are pushed away from regions of high density, whereas they are attracted to high-density regions for negative $A$. Note that the normalization of $P_{\mathrm{int}}$ is proportional to time
$\int_{-\infty}^{\infty}dxP_{\mathrm{int}}(x,t_{n})=n~,$ (10)
@@ II.3 Simulation details
To implement the mean-density interaction, we consider ensembles of up to $N=128$ particles. The mean integrated density $P_{\mathrm{int}}(x,t_{n})$ is collected as a histogram with a narrow bin width $\Delta x=0.1$. To achieve a smooth mean integrated density for our moderately large ensemble sizes, we replace the $\delta$ functions in the definition (5) of $P_{\mathrm{int}}(x,t_{n})$ by Gaussians of variance 0.25. We have confirmed that small changes of the histogram bin width and the variance of the smoothing Gaussian only lead to minuscule changes of the results  [^2] If the $\delta$ functions in the definition (5) of $P_{\mathrm{int}}(x,t_{n})$ are not smoothed, however, the histogram of the mean integrated density $P_{\mathrm{int}}(x,t_{n})$ develops an unphysical oscillatory instability.. The gradient in the definition of the force (9) is computed via a simple two-point formula directly from the histogram. To ensure the robustness of the results, we have also performed test calculations using a higher-order (four-point) gradient formula as well as calculations that do not use a histogram at all but analytically compute the derivative of the sum of the accumulated Gaussians. The latter method is restricted to times $t\lessapprox 10^{4}$ because of the high numerical effort of keeping track of all the Gaussians. All gradient algorithms give the same results (within our error bars) for the available simulation times.
@@ III Results: Mean-squared displacement
Figure 2: Mean-squared displacement $\langle x^{2}\rangle$ of FBM with mean-density interaction vs. time $t$ for $\alpha=0.7$ and several values of the interaction strength AA. The data are averages over 60 ensembles of 128 random walkers each. The resulting statistical errors are much smaller than the symbol size. The dashed line represents a fit of the long-time behavior for $A=2/5$ with $\langle x^{2}\rangle=c\,t^{4/3}$ whereas the dash-dotted line shows the FBM relation $\langle x^{2}\rangle=\sigma^{2}t^{0.7}$. Inset: crossover time $\tau_{x}$ from FBM to interaction-dominated behavior vs. interaction strength AA. The solid line is a power law fit, yielding an exponent of 1.05-1.05.
@@
The trajectories of non-interacting FBM with $\alpha=0.7$ shown in panel (b) are highly jittery (antipersistent) because of the negative correlations of subdiffusive ($\alpha<1$) fractional Gaussian noise. This jittery motion is still visible in the presence of the mean-density interaction shown in panel (a). However, it is superposed onto a more regular motion away from the starting point (clearly visible for the trajectories furthest away from the origin (in the tails of the probability density). The comparison with panel (c) demonstrates that the trajectories of FBM with mean-density interaction do not resemble those of non-interacting FBM for $\alpha=4/3$ even though the mean-squared displacements of both processes evolve as $\langle x^{2}\rangle\sim t^{4/3}$. The $\alpha=4/3$ FBM trajectories are less jittery, even on short time scales because the fractional Gaussian noise for $\alpha>1$ is positively correlated  [^3] For non-interacting FBM in $d$ dimensions, the probability of trajectory crossings can be estimated from scaling arguments. During time $t$, a trajectory explores a length proportional to $t^{\alpha/2}$, yielding a coincidence rate proportional to $t^{-d\alpha/2}$. The number of crossings between times 0 and $t$ then reads $N(t)\sim\int_{0}^{t}d\tau\,\tau^{-d\alpha/2}$. Crossings proliferate for $d\alpha/2<1$ which always holds in $d=1$. For the parameters relevant for the serotonergic fiber problem, $d\alpha/2>1$, and crossings are rare in free space. .
@@ IV Scaling theory
If we assume, however, that the length scale $b(t)$ increases more slowly than $t^{2/3}$, the hypothetical contribution of the forces to the displacement would increase faster than $t^{2/3}$, leading to a contradiction. The scaling theory therefore predicts that one-dimensional FBM with mean-density interaction in free space is dominated by the fractional Gaussian noise (and behaves like regular FBM) for $\alpha>4/3$ (i.e., $\gamma<2/3)$, whereas it is interaction-dominated for $\alpha<4/3$ (i.e., $\gamma>2/3$). This yields the following mean-squared displacement behaviors,
$\langle x^{2}\rangle\sim\left\{\begin{array}[]{ll}t^{4/3}&\qquad\textrm{for }\alpha<4/3~,\\ t^{\alpha}=t^{2-\gamma}&\qquad\textrm{for }\alpha>4/3~.\end{array}\right$. (19)
@@
$\tau_{x}\sim\left(\frac{A}{\sigma}\right)^{2/(3\alpha-4)}~.$ (20)
This relation holds in the interaction-dominated regime, $\alpha<4/3$. For the case of $\alpha=0.7$, the exponent in the power law (20) evaluates to 1.053-1.053, in excellent agreement with the data in the inset of Fig. 2.
@@ V Results: Probability densities
where $\Gamma$ is the upper incomplete $\Gamma$ function (for details, see Appendix B). The function (24) has a maximum (with a cusp) at $x=0$ and a Gaussian tail (up to a power-law prefactor) for large $x$.
@@
$P(x,t)=\frac{1}{b(t)}Z\left[\frac{x}{b(t)}\right]~,$ (25)
@@
Despite its highly non-Gaussian form, the probability density in the interaction-dominated regime fulfills the scaling form (25) with $b(t)=x_{\mathrm{rms}}(t)=\sqrt{\langle x^{2}(t)\rangle}$, as can be seen in Fig. 7(b). The small deviations from perfect scaling collapse for the shortest time in the figure can be attributed to finite-time effects. Specifically, the force terms do not completely dominate at $t=2^{18}$, and the fractional Gaussian noise produces the small tails at large $|x|$.
@@
VI Finite-$N$ fluctuations
So far, we have considered large particle numbers ($N\gg 1$). In the limit $N\to\infty$, the integrated density of a given ensemble of $N$ particles becomes identical to its average over all noise realizations, i.e., finite-$N$ fluctuations of $P_{\mathrm{int}}$ and the resulting forces are suppressed. In this section, we study the effects of a finite particle number (ensemble size) $N$. We focus on the interaction-dominated regime $\alpha<4/3$ because the forces do not affect the asymptotic behavior in the noise dominated regime $\alpha>4/3$.
@@ VI Finite-$N$ fluctuations
Figure 8: Mean-squared displacement $\langle x^{2}\rangle$ of FBM with mean-density interaction vs. time $t$ for $\alpha=0.7$, interaction strength $A=1/40$ and different particle numbers NN. The data are averages over 2048 ensembles ($N=1$), 1024 ensembles ($N=2$), 128 ensembles ($N=16$), and 16 ensembles ($N=128$). The dashed lines are fits of the long-time behavior for $N=1$ and $N=128$ with $\langle x^{2}\rangle=c\,t^{4/3}$.
@@
Figure 9: Integrated density $P_{\mathrm{int}}(x,t)$ at time $t=2^{20}$ for $\alpha=0.7$, interaction strength $A=1/40$ and different particle numbers NN. The data are averages over 4096 ensembles ($N=1$), 2048 ensembles ($N=2$), 256 ensembles ($N=16$), and 32 ensembles ($N=128$).
@@ VII Generalizations
Repeating the arguments at the end of Sec.  IV, we conclude that the motion of FBM with mean-density interaction will be interaction dominated if the FBM anomalous diffusion exponent $\alpha$ is smaller than $2\delta$. In this case the mean-squared displacement is expected to behave as $\langle\mathbf{x}^{2}\rangle\sim t^{2\delta}$. If, however, $\alpha>2\delta$, then the motion will be dominated by the fractional Gaussian noise leading to $\langle\mathbf{x}^{2}\rangle\sim t^{\alpha}$. Equation (28) shows that $\delta$ decreases with increasing space dimensionality $d$. This implies that the marginal value of $\alpha$, below which the interactions dominate over the noise, decreases with increasing $d$. The result that the interaction effects are strongest in one dimension and decrease with increasing $d$ is perhaps not unexpected as crowding is more easily achieved in lower dimensions.
Let us specifically consider the cases $d=2$ and $d=3$ for linear gradient forces, $\lambda=1$. In two dimensions, Eq. (28) reduces to $\delta=1/2$. This implies that $\alpha=2\delta=1$ is the marginal value of $\alpha$ that separates the noise-dominated and interaction-dominated regimes in two dimensions. We note that this agrees with the finding that $d=2$ is the critical dimension of the so-called “true” or myopic self-avoiding random walk [38] discussed in Sec.  VIII. Analogously, for $d=3$, we find $\delta=2/5$ from Eq. (28), implying a critical $\alpha=0.8$.
@@ Acknowledgements.
Data availability
The data that support the findings of this article are openly available [54].
@@ Appendix A Interaction with finite decay time
$P_{\mathrm{tot}}(x,t_{n})$ $=$ $\sum_{m=1}^{n}e^{-(t_{n}-t_{m})/\tau_{d}}\,\sum_{j=1}^{N}\delta[x-x^{(j)}_{m}$$\quad$ (29)
@@
To ensure that the three time regimes predicted by the scaling theory (see Table 1) are well separated, we have chosen a comparatively strong interaction, $A=1/5$, for these simulations. This reduces the crossover time $\tau_{x}$ from the initial FBM regime to the interaction-dominated regime. The data for $\alpha=0.4$ (shown in the upper panel of the figure) clearly feature the three predicted regimes. Initially, for time $t<\tau_{x}\approx 10$, the mean-squared displacement follows the FBM relation $\langle x^{2}\rangle\sim t^{0.4}$. At $\tau_{x}$, it crosses over to the interaction-dominated behavior $\langle x^{2}\rangle\sim t^{4/3}$ established in the main part of the paper. After the second crossover at $t=\tau_{d}$, the long-time behavior reaches the second interaction-dominated regime with $\langle x^{2}\rangle\sim t^{2/3}$. This can be seen clearly in the figure where the solid line for $\tau_{d}=10^{3}$ represents a power-law fit with the exponent $2/3$.
@@ Appendix B Integrated density of FBM
$P_{\mathrm{int}}(x,t)$ $=$ $\int_{0}^{t}d\tau P(x,\tau)$ (32)
$=$ $\int_{0}^{t}d\tau\frac{1}{\sqrt{2\pi\sigma^{2}\tau^{\alpha}}}\exp{\left(-\frac{x^{2}}{2\sigma^{2}\tau^{\alpha}}\right)}~.$
@@
$P_{\mathrm{int}}(x,t)=\frac{|x|^{2/\alpha-1}}{\alpha\pi^{1/2}(2\sigma^{2})^{1/\alpha}}\int_{x^{2}/(2\sigma^{2}t^{\alpha})}^{\infty}dz\,z^{-1/\alpha-1/2}e^{-z}~.$ (33)
The integral over $z$ yields the upper incomplete $\Gamma$ function, which concludes the derivation of Eq. (24). It is worth emphasizing that expression (24) fulfills the scaling form (11) with $b(t)=x_{\mathrm{rms}}(t)=\sigma t^{\alpha/2}$. This can be seen explicitly by rewriting (24) as
$P_{\mathrm{int}}(x,t)$ $=$ $\frac{t}{\sigma t^{\alpha/2}}\,Y\left(\frac{x}{\sigma t^{\alpha/2}}\right)~,$ (34)